OpenStudy (anonymous):

write a rule for the sequence -4,5,20,41,68 write a rule for the sequence -4,5,20,41,68 @Mathematics

OpenStudy (anonymous):

-4, 5 , 20, 41, 68 -4 +9 =5 5+9+6 =20 20+15+6 =41 41+21+6=68

OpenStudy (precal):

It is not arithmetic nor geometric. I wonder what it could be?

OpenStudy (anonymous):

i'm not sure. i'm also supposed to use sigma notation to find the sum of the first 10 terms

OpenStudy (precal):

usually the sequences are arithmetic or geometric, but a sequence can be anything at all.